A thin film of acetone (n = 1.25) coats a thick glass plate (n = 1.50). White light is incident normal to the film. In the reflections, fully destructive interference occurs at 510 nm and fully constructive interference at 574 nm. Calculate the thickness of the acetone film.
Answer #1

we know that

2L = (m + 0.5)lambda/n2

For m = 0

L₁ = (0 · 574) / (2 · 1.25) = 0 nm

L₂ = (0 + 0.5) · [510 ÷ (2 · 1.25)] = 102 nm

For m = 1

L₁ = (1 · 574) / (2 · 1.25) = 229.6 nm

L₂ = (1 + 0.5) · [510 ÷ (2 · 1.25)] = 306 nm

For m = 2

L₁ = (2 · 574) / (2 · 1.25) = 459.2 nm

L₂ = (2 + 0.5) · [510 ÷ (2 · 1.25)] = 510 nm

For m = 3

L₁ = (3 · 574) / (2 · 1.25) = 688.8 nm

L₂ = (3 + 0.5) · [510 ÷ (2 · 1.25)] = 714 nm

For m = 4

L₁ = (4 · 574) / (2 · 1.25) = 918 nm

L₂ = (4 + 0.5) · [510 ÷ (2 · 1.25)] = 918 nm


Hence, the thickness of the acetone film is L = 918 nm